# An Approach for Systematic Decomposition of Complex LLM Tasks

### v2

@@ Abstract
Large Language Models (LLMs) suffer from reliability issues on complex tasks, as existing decomposition methods are heuristic and rely on agent or manual decomposition. This work introduces a novel, systematic decomposition framework that we call Analysis of CONstraint-Induced Complexity (ACONIC), which models the task as a constraint problem and leveraging formal complexity measures to guide decomposition. On combinatorial (SATBench) and LLM database querying tasks (Spider), we find that by decomposing the tasks following the measure of complexity, agent can perform considerably better (10-40 percentage point).
@@ 1 Introduction
A growing body of work addresses this limitation through task decomposition. Instead of solving a task monolithically, these methods break it into smaller, more tractable subtasks. One popular line of work, starting frorm chain-of-though Wei et al. (2022), is to use LLMs for decomposition Yao et al. (2023); Khot et al. (2023); Pourreza and Rafiei (2023); Chen et al. (2024). Other methods rely on domain experts to decompose the task into workflows and provide access to tools that shoulder parts of the task (Wang et al., 2024; Singh et al., 2024). We summarize these methods as the grey path in Figure˜1.
@@
We use SAT-Bench (Wei et al., 2025b) and the NL2SQL Spider (Yu et al., 2019) benchmark to study this framework. We find that the task measures define frontiers of difficulty that separate tasks easily solvable by LLMs from those that are nearly impossible without structural assistance. In addition, decomposition based on complexity improves task accuracy by $9-40\%$ as compared to chain-of-thought decomposition.
2 Reduction to Constraint Representation
At a high level, our reduction happens in two steps. The first models the agent’s task using a state-based framework that captures the interaction between the agent and its observations of the system environment. This enables us to model the agent’s planning as a satisfiability (PaS) problem Selman et al. , and further reduce this to a standard constraint satisfiability problem. In our experiments, we develop automatic reduction procedures for two benchmarks.
2.1 Reduction to Planning Problems
The state based framework of agent operations (e.g. Agent Path Finding  Surynek et al. (2016)) can be formulated as PaS problems Selman et al. . The framework transfers world states and actions into finite state abstractions
$P=\langle F,A,I,G\rangle,$
where $F$ is a finite set of propositional fluents describing world facts, $A$ is a finite set of actions, and $I$, $G$ determines the initial and goal fluents. For a given action $a$, $\mathcal{P}(a)\subseteq F$ determines the preconditions, $\mathcal{A}(a)\subseteq F$ determines the fluents that become true, and $\mathcal{D}(a)\subseteq F$ determines the fluents that become false. An agent task becomes determining a sequence of actions $\langle a_{1},\dots,a_{k}\rangle$ such that, starting from $I$, each $a_{i}$ is applicable (its preconditions hold), and executing its effects (adding and deleting fluents) transfers current state to a new state. The plan is valid if the final state after executing all actions satisfies all goals in $G$.
Consider a toy meeting scheduling problem:
@@ 2.1 Reduction to Planning Problems
In the toy example above, let us define $x_{p}^{(t,\ell)}$ to be a Boolean action variable indicating that person $p$ is invited to location $\ell$ at time $t$, where $\mathcal{P}$, $\mathcal{T}$, and $\mathcal{L}$ denote the global set of people, times, and locations, respectively. For each person $p$, let $\mathcal{A}_{p}\subseteq\mathcal{T}\times\mathcal{L}$ be the set of feasible (time, location) pairs. For each pair $(p,p^{\prime})$, define the pairwise feasible set $\mathcal{F}_{p,p^{\prime}}=\mathcal{A}_{p}\cap\mathcal{A}_{p^{\prime}}$ as given in the system record.
We model the state as the set of value assignments $\{x_{p}^{(t,\ell)}\mid p\in\mathcal{P},(t,\ell)\in\mathcal{T}\times\mathcal{L},x_{p}^{(t,\ell)}\in\{0,1\}\}$, indicates whether participant $p$ attends slot $(t,\ell)$. An action consists of assigning some $x_{p}^{(t,\ell)}=1$.Then a target set can be constructed by requiring that for each desired meeting pair $(p,p^{\prime})\in\{(A,B),(A,C)\}$, exactly one common feasible slot is chosen where both $p$ and $p^{\prime}$ are present. Formally, defining meeting at $\ell,t$ for $(p,p^{\prime})$ to be $M_{p,p^{\prime}}^{(t,\ell)}:=x_{p}^{(t,\ell)}\wedge x_{p^{\prime}}^{(t,\ell)}$, the goal condition can be written as:
$G:\;$ $\bigwedge_{(p,p^{\prime})}\operatorname{\mathrm{EO}}\!\big\{M_{p,p^{\prime}}^{(t,\ell)}\mid(t,\ell)\in\mathcal{F}_{p,p^{\prime}}\big\}\wedge\;$
$\bigwedge_{(t,\ell)}\neg\big(M_{A,C}^{(t,\ell)}\wedge M_{A,B}^{(t,\ell)}\big),$ (1)
where $\mathrm{EO}$ denotes Exactly One, which can be easily represented with at least one, and at most one. All $x_{p}^{(t,\ell)}$ initialized to be False.
The agent must produce a sequence of actions that comply with the preconditions and effects defined in $P=\langle F,A,I,G\rangle$, thus transiting the states toward a goal-satisfying state.
Our objective is to find a problem decomposition strategy that provide observations over the subset of the problem. It should ensure that the agent’s local consistency over observed subgraphs guarantees global satisfiability of $G$.
2.2 Reduction to Constraint Satisfaction
Given the planning-as-satisfaction formulation $P=\langle F,A,I,G\rangle$, we can reduce the problem to a constraint satisfaction instance by encoding the preconditions $f_{p}\in\mathcal{P}(a)$, add effects $f_{a}\in\mathcal{A}(a)$, and delete effects $f_{d}\in\mathcal{D}(a)$ of each action as Boolean dependent constraints between fluents and actions. Additionally, we include a $k$-wise constraint involving all variables in $\mathcal{P}(a)\cup\mathcal{A}(a)\cup\mathcal{D}(a)$ to enforce frame consistency and prevent anomalous models Selman et al. where fluents change truth values without any triggering action. By additionally encoding the goal states $G$ as constraints, solving the resulting CSP is equivalent to solving the original PaS problem.
In constraint processing theory, tree decomposition Bodlaender (1998) on a constraint graph provides desirable guarantees of consistency and modularity: when traversing along the tree, any variable assignment made within one subgraph does not affect previously satisfied clauses except through shared boundary variables (which is always included in current subproblem). We therefore chose tree decomposition to separate the CSP into subproblems, the completion of which lead to global success of the problem. This ensures that local consistency over the bags implies global consistency of the full CSP solution.
Bodlaender (1998) show existence of a tree decomposition $\mathcal{D}^{*}$ with the smallest maximum bag size (treewidth), which characterizes the intrinsic problem complexity. Leveraging this prior work, ACONIC uses minimal tree decomposition to partition reasoning tasks into the minimal locally consistent subtasks while preserving global satisfiability.
@@ 3 Experiments
We now evaluate our framework against chain-of-thought decomposition Wei et al. (2022) on two benchmarks: SAT-based story problems (SAT-bench) and natural language to SQL (Spider).
@@ 3.1 SAT-Bench
SAT-Bench Wei et al. (2025a) uses (i) an underlying SAT problem to construct (ii) a natural-language story describing the same constraints, and (iii) an alignment from story entities to their SAT representations. Instead of asking the LLM for an immediate satisfiability judgment, we prompt it to produce value assignments step-by-step to ensure a complete understanding of the assignment process. In this process, the state–action dependency is modeled by casting SAT as a PaS problem, where the goal state is defined by the original SAT-Bench CNF representation.
@@
We evaluated Llama-3-70B on all tasks and Claude3.5-Sonnet on half the tasks (randomly sampled). Figure˜2 plots the success/failure of each task by its complexity measures, and exhibits frontiers of difficulty beyond which tasks are too difficult or too simple. For the baseline, there appears to be a fixed “total task complexity” evidenced by the trade-off between problem treewidth and number of bags (left column). In contrast, ACONIC decomposition pushes the frontier outward and is able to successfully complete more complex tasks. Overall, ACONIC increases the task completion rate from $49.3\%\to 58.1\%$ using Claude, and $21.5\%\to 36.5\%$ using LLaMA—a $9-15\%$ improvement on both models.
@@ 3.2 Natural Language to SQL
Natural language to SQL (NL2SQL) tasks translate a natural language query over a database into a desired SQL query. A major challenge is constructing the join condition to connect tables referenced in the query. For example, the schema in Figure˜3 contains 4 tables about singers and concerts. Although the NL query only refers to singers and concerts, the LLM needs to infer that singer_in_concern is needed for a valid join path.
@@
We use the popular Spider NL2SQL benchmark dataset Yu et al. (2019). It contains hundreds of databases, each containing up to 37 tables, 90 foreign keys, and potentially over 100 columns. Each task consists of a database schema $S$, the natural language (NL) query $q$, and a ground-truth SQL query $q^{*}$. The original benchmark submits $(S,q)$ to the LLM, which can result in incorrect or invalid join conditions. Given the tables referenced in $q$, ACONIC decomposes the schema into subgraphs with minimized maximal complexitieså and constructs a workflow (grey arrows). For instance, the first subtask gives the agent $q$ and the schemas for subgraph singer-singer_concern (red arrow), and asks it to construct the appropriate WITH clause. The next subtask does the same with the second subgraph (orange arrow), but is also provided the output from the previous subtask. A verification agent cleans up minor errors to build the final SQL query.
Figure 4: ACONIC result on Llama3-70B
Figure 5: Baseline CoT on Llama3-70B
Figure˜4 and Figure˜5 respectively report accuracy for ACONIC and the CoT baseline. Each bar reports the best of 1, 2, or 3 agent trajectories. We find that while increasing the number of trajectories slightly improves the accuracy, decomposition using ACONIC increases the accuracy over CoT by $\approx 30-40\%$ across the conditions.
@@ 4 Conclusions
ACONIC is a principled framework for decomposing complex LLM tasks by reducing them to constraint satisfaction problems and quantifying complexity via treewidth. Unlike heuristic methods, ACONIC minimizes local complexity while preserving global satisfiability. Across SATBench and Spider, this yielded frontiers of task difficulty that define LLM reasoning limits. Complexity-based decomposition improved completion rates by up to $15\%$ on SAT-bench and up to $40\%$ on SPIDER. These results suggest a path towards theoretically grounded, reliable multi-step LLM systems.
@@ 5 Limitations
The tasks that we evaluated can be conveniently modeled as constraint satisfiability problems. Although many other practical problems, such as deadlock detection in databases, resource scheduling in operating systems, task placement in distributed systems, and control/data-flow construction in agentic programming, can be formulated as constraint satisfiability problems, they often cannot be logically represented completely, either due to question ambiguity, lack of transparancy of the agent actions, or fuzzy contextual information. In these cases, future work might study hybrid decomposition approaches that mix logical and common-sense constraints in a task.
@@ A.2 CSP Encoding of SAT Clauses
We build a CSP instance $\mathcal{C}=\langle X,D,C\rangle$ as follows. $X=\{x_{1},\dots,x_{n}\}$ (one variable per SAT variable $v_{i}$) with domain of each variable being D=0,1D={0,1}
@@
We encode the variable $V$ to CSP variables with initiation of 0 (False) and the clauses $C_{j}=\bigl(\bigvee_{r=1}^{k_{j}}\ell_{j,r}\bigr)$ with literal $\ell_{j,r}\in\{v_{r},\neg v_{r}\}$ as i wise constraint $\mathbf{T}_{j}$ including all the literals in $C_{j}$.
@@ A.3 Tree Decomposition on the Induced CSP Graph
To analyze structural complexity, we perform tree decomposition on this graph, which partitions the variable set into overlapping subgraphs (bags) while minimizing the size of the largest bag. The minimum achievable largest bag size minus one corresponds to the graph’s treewidth, which we use as a key measure of problem complexity and locality in constraint reasoning Bodlaender (1998) in chart 2(a), 2(b), 2(c), 2(d)
@@ B.4 Experiment Workflow
We compare the baseline Chain-of-Thought (CoT) decomposition with the proposed ACONIC framework on the NL2SQL task. In both settings, the agent interacts with the environment over multiple rounds, constructing the final SQL query.
@@
In the CoT baseline, the agent receives the entire database schema at once and is asked to generate the complete SQL query that satisfies the NLQ. This configuration represents global reasoning over the full relational structure, without any decomposition or locality constraints.

### v3

@@ Abstract
Large Language Models (LLMs) suffer from reliability issues on complex tasks, as existing decomposition methods are heuristic and rely on agent or manual decomposition. This work introduces a novel, systematic decomposition framework that we call Analysis of CONstraint-Induced Complexity (ACONIC), which models the task as a constraint problem and leverages formal complexity measures to guide decomposition. On combinatorial (SATBench) and LLM database querying tasks (Spider), we find that by decomposing the tasks following the measure of complexity, agent can perform considerably better.
An Approach for Systematic Decomposition of Complex LLM Tasks
Tianle Zhou§, Jiakai Xu§, Guanhong Liu§, Jiaxiang Liu§, Haonan Wang§, Eugene Wu∗§ §Columbia University
@@ 1 Introduction
A growing body of work addresses this limitation through task decomposition. Instead of solving a task monolithically, these methods break it into smaller, more tractable subtasks. One popular line of work, starting from chain-of-thought Wei et al. (2022), is to use LLMs for decomposition Yao et al. (2023); Khot et al. (2023); Pourreza and Rafiei (2023); Chen et al. (2024). Other methods rely on domain experts to decompose the task into workflows and provide access to tools that shoulder parts of the task (Wang et al., 2024; Singh et al., 2024). We summarize these methods as the grey path in Figure˜1.
@@
We use SAT-Bench (Wei et al., 2025) and the NL2SQL Spider (Yu et al., 2019) benchmark to study this framework. We find that the task measures define frontiers of difficulty that separate tasks easily solvable by LLMs from those that are nearly impossible without structural assistance. In addition, complexity-guided decomposition improves completion on SATBench and outperforms Tree-of-Thoughts on Spider by 3–8 accuracy points across difficulty levels.
2 Method
At a high level, we find that a class of agent tasks that admit constraint formulations can be reduced to constraint satisfaction problems (CSPs), such that solving the CSP is equivalent to solving the original task. This formulation makes explicit the dependency structure among task variables, which allows us to leverage tree decomposition from constraint processing theory to decompose complex tasks into smaller, weakly coupled subproblems that can be reasoned about locally and composed consistently.
2.1 Reduction to Constraint Satisfaction
We model a class of agent tasks as instances of constraint satisfaction problems (CSPs). A CSP is defined by a tuple $\mathcal{C}=\langle X,D,\mathcal{R}\rangle$, where $X$ is a finite set of variables, $D$ assigns each variable a finite domain, and $\mathcal{R}$ is a set of constraints over subsets of variables.
Within this framework, the state of the agent is a (partial) assignment to variables in $X$, and an action corresponds to proposing or updating values for a subset of variables, or acquiring information that refines the constraints. The agent’s objective is to construct a complete assignment that satisfies all constraints in $\mathcal{R}$.
We illustrate this abstraction using a simple meeting scheduling example.
@@ 2.1 Reduction to Constraint Satisfaction
In this example, resolving conflicts is non-trivial because assignments for one meeting constrain the feasible choices for the other. For instance, choosing a time for Alice and Charlie may invalidate the only feasible option for Alice and Bob.
We introduce Boolean variables $x_{p}^{(t,\ell)}$ indicating whether person $p$ attends location $\ell$ at time $t$. Let $\mathcal{P}$ denote the set of participants, $\mathcal{T}$ time slots, and $\mathcal{L}$ locations. For each person $p$, let $\mathcal{A}_{p}\subseteq\mathcal{T}\times\mathcal{L}$ denote feasible slots, and for each pair $(p,p^{\prime})$ let $\mathcal{F}_{p,p^{\prime}}=\mathcal{A}_{p}\cap\mathcal{A}_{p^{\prime}}$.
The goal is to select exactly one feasible meeting slot for each desired pair $\mathcal{F}_{A,B}$ and $\mathcal{F}_{A,C}$, and to ensure that no one attend two meetings simultaneously. This leads to a set of constraints over the variables $\{x_{p}^{(t,\ell)}\}$ (As demonstrated in appendix E).
The agent must produce a valid initiation of the variables $X$ that satisfies all the constraints. Our objective is to find a problem decomposition strategy that provide the agent with observations over a subset of the problem.
2.2 Tree Decomposition for Agent Tasks
Figure 2: CSP graph (and its tree decomposition) for Meeting Scheduling example. Constructed based on CSP representation in appendix E
As illustrated in 2, the CSP induces a constraint graph $G=(X,E)$, where variables are nodes and an edge connects two variables if they co-occur in some constraint. It provides a structural abstraction of the task, capturing which task variables are mutually dependent through constraints. This representation allows us to reason about task structure independently of concrete values, and motivates decomposition strategies that group tightly coupled variables into localized subproblems. We therefore explore how tree decomposition Bodlaender (1998) of this graph can be used to organize and simplify agent reasoning.
Tree decomposition maps $G$ into a tree of bags $\mathcal{D}=(T,\{B_{i}\})$, where each bag $B_{i}\subseteq X$ satisfies:
  • Every constraint is contained in some bag.
  • For any variable $x$, the set of bags containing $x$ forms a connected subtree.
Following these properties, tree decomposition ensures that if each bag enforces local consistency on its induced constraints, then these local solutions can be merged into a global satisfying assignment as long as shared variables agree along the tree.
Among all tree decompositions $\mathcal{D}=(T,\{B_{i}\})$ of $G$, there exists an optimal decomposition $\mathcal{D}^{*}\in\arg\min_{\mathcal{D}}\ \max_{i}|B_{i}|.$ The corresponding minimum possible maximum bag size, $b^{*}(G):=\min_{\mathcal{D}}\max_{i}|B_{i}|,$ induces the treewidth of $G$ as $\mathrm{tw}(G)=b^{*}(G)-1$. We use this quantity to characterize the intrinsic structural complexity of the underlying constraint problem: smaller $\mathrm{tw}(G)$ implies that the task can be decomposed into smaller, more weakly-coupled subproblems.
We exploit this structure to decompose reasoning into minimal, locally consistent subtasks. Each bag corresponds to a subproblem over a small variable set. The agent observes and reasons following the subproblems, while keeping consistency over intersections and ensuring global satisfiability. This leads to a principled decomposition strategy whose complexity is governed by the intrinsic treewidth of the task.
@@ 3 Experiments
We now evaluate our framework against chain-of-thought decomposition Wei et al. (2022) and Tree of Thoughts Yao et al. (2023) on two datasets: SAT-based story problems (SAT-bench) and natural language to SQL (Spider).
@@ 3.1 SAT-Bench
SAT-Bench Wei et al. (2025) uses (i) an underlying SAT problem to construct (ii) a natural-language story describing the same constraints, and (iii) an alignment from story entities to their SAT representations. Instead of asking the LLM for an immediate satisfiability judgment, we prompt it to produce value assignments step-by-step to ensure a complete understanding of the assignment process. In this process, the agent’s state is a partial assignment to SAT variables, and progress corresponds to extending this assignment while maintaining consistency with the underlying constraints. Concretely, we treat the SAT-Bench CNF as a Boolean CSP, where the goal is to find an assignment that satisfies all clauses.
@@
We evaluated Llama-3-70B on all tasks and Claude3.5-Sonnet on half the tasks (randomly sampled). Figure˜3 plots the success/failure of each task by its complexity measures, and exhibits frontiers of difficulty beyond which tasks are too difficult or too simple. For the baseline, there appears to be a fixed “total task complexity” evidenced by the trade-off between problem treewidth and number of bags (left column). In contrast, ACONIC decomposition pushes the frontier outward and is able to successfully complete more complex tasks. Overall, ACONIC increases the task completion rate from $49.3\%\to 58.1\%$ using Claude, and $21.5\%\to 36.5\%$ using LLaMA—a $9-15\%$ improvement on both models.
@@ 3.2 Natural Language to SQL
Natural language to SQL (NL2SQL) tasks translate a natural language query over a database into a desired SQL query. A major challenge is constructing the join condition to connect tables referenced in the query. For example, the schema in Figure˜4 contains 4 tables about singers and concerts. Although the NL query only refers to singers and concerts, the LLM needs to infer that singer_in_concert is needed for a valid join path.
@@
We use the popular Spider NL2SQL benchmark dataset Yu et al. (2019). It contains hundreds of databases, each containing up to 37 tables, 90 foreign keys, and potentially over 100 columns. Each task consists of a database schema $S$, the natural language (NL) query $q$, and a ground-truth SQL query $q^{*}$. The original benchmark submits $(S,q)$ to the LLM, which can result in incorrect or invalid join conditions. Given the tables referenced in $q$, ACONIC decomposes the schema into subgraphs with minimized maximal complexities and constructs a workflow (grey arrows). For instance, the first subtask gives the agent $q$ and the schemas for subgraph singer-singer_concert (red arrow), and asks it to construct the appropriate WITH clause. The next subtask does the same with the second subgraph (orange arrow), but is also provided the output from the previous subtask. A verification agent cleans up minor errors to build the final SQL query.
Figure 5: Accuracy performance of Tree of Thought (ToT) and ACONIC (Ours) on Spider EASY, MEDIUM, HARD, and EXTRA tasks
@@ 4 Conclusions
ACONIC is a principled framework for decomposing complex LLM tasks by reducing them to constraint satisfaction problems and quantifying complexity via treewidth. Unlike heuristic methods, ACONIC minimizes local complexity while preserving global satisfiability. Across SATBench and Spider, this yielded frontiers of task difficulty that define LLM reasoning limits. Complexity-based decomposition improved completion rates by up to $15\%$ on SAT-bench. On Spider datasets, the accuracy was increased over ToT by up to 7.5 points (and +5.4 points overall). These results suggest a path towards theoretically grounded, reliable multi-step LLM systems. Figure˜5 reports Spider accuracy for ACONIC and Tree-of-Thoughts (ToT). Across different difficulty levels, ACONIC consistently outperforms ToT by 3.6–7.5 accuracy points, with an overall gain of +5.4 points (50.68% $\rightarrow$ 56.09%).
@@ 5 Limitations
The tasks that we evaluated can be conveniently modeled as constraint satisfiability problems. Although many other practical problems, such as deadlock detection in databases, resource scheduling in operating systems, task placement in distributed systems, and control/data-flow construction in agentic programming, can be formulated as constraint satisfiability problems, they often cannot be logically represented completely, either due to question ambiguity, lack of transparency of the agent actions, or fuzzy contextual information. In these cases, future work might study hybrid decomposition approaches that mix logical and common-sense constraints in a task.
@@ A.2 CSP Encoding of SAT Clauses
We build a CSP instance $\mathcal{R}=\langle X,D,C\rangle$ as follows. $X=\{x_{1},\dots,x_{n}\}$ (one variable per SAT variable $v_{i}$) with domain of each variable being $D_{x_{i}}=\{0,1\}$
@@
We encode the variable $V$ to CSP variables with initiation of 0 (False) and the clauses $C_{j}=\bigl(\bigvee_{r=1}^{k_{j}}\ell_{j,r}\bigr)$ with literal $\ell_{j,r}\in\{v_{r},\neg v_{r}\}$ as clause-wise constraint $\mathbf{T}_{j}$ including all the literals in $C_{j}$.
@@ A.3 Tree Decomposition on the Induced CSP Graph
To analyze structural complexity, we perform tree decomposition on this graph, which partitions the variable set into overlapping subgraphs (bags) while minimizing the size of the largest bag. The minimum achievable largest bag size minus one corresponds to the graph’s treewidth, which we use as a key measure of problem complexity and locality in constraint reasoning Bodlaender (1998) in chart 6(a), 6(b), 6(c), 3(d)
@@ A.4 Experiment Workflow
Algorithm 1 SAT-Bench Workflow
1:Input: NL story $Q$, CNF clauses $\Phi$, mapping $\psi$
2:Build CSP:
3:   $X\leftarrow$ variables in $\Phi$
4:for each clause $C$ in $\Phi$ do
5:    $T_{C}\leftarrow$ all satisfying assignments of variables in $C$
6:end for
7:Build constraint graph:
8:for each clause $C$ in $\Phi$ do
9:    connect variables in $C$ in graph
10:end for
11:Tree decomposition:
12:   Bags $\leftarrow$ tree_decompose(graph)
13:Agent interaction:
14:   State $\leftarrow$ empty assignment
15:while not solved and budget not exceeded do
16:  if CoT / ToT / DaC then
17:     Obs $\leftarrow$ (current violation, all clauses in $\Phi$)
18:  else if ACONIC then
19:     Violated $\leftarrow$ clauses violated under State
20:     Bag $\leftarrow$ bag containing variables of Violated
21:     Obs $\leftarrow$ (current violation, clauses inside Bag)
22:  end if
23:  Action $\leftarrow$ LLM(State, $Q$, $\psi$, Obs)
24:  State $\leftarrow$ update assignment(Action)
25:end while
26:Return final assignment
@@
A.5 Implementation
We implement SAT-Bench experiment as a Terminal-Bench task Team (2025) to ensure reproducibility and comparability between agents.
@@ B.4 Experiment Workflow
We compare the baseline Tree-of-Thought (ToT) heuristic decomposition method with the proposed ACONIC framework on the NL2SQL task. In both settings, the agent interacts with the environment over multiple rounds, constructing the final SQL query.
Algorithm 2 Spider Workflow
1:Input: NL query NLQ, database schema $S$
2:Build CSP:
3:   $X\leftarrow$ tables in $S$
4:Build constraint graph (foreign keys only):
5:for each foreign key constraint do
6:    connect corresponding tables in graph
7:end for
8:Tree decomposition:
9:   Bags $\leftarrow$ tree_decompose(graph)
10:Agent interaction:
11:   CTEs $\leftarrow$ empty list
12:while unresolved tables remain do
13:  if ToT then
14:     Obs $\leftarrow$ (current SQL, full schema $S$)
15:  else if ACONIC then
16:     Target $\leftarrow$ tables referenced in NLQ
17:     Bag $\leftarrow$ bag containing unresolved Target tables
18:     Obs $\leftarrow$ (CTEs, schema restricted to Bag)
19:  end if
20:  Fragment $\leftarrow$ LLM(NLQ, Obs, Schema $S$)
21:  append Fragment to CTEs
22:end while
23:LLM merging:
24:   SQL $\leftarrow$ LLM_based_merge_and_resolve(CTEs)
25:Return final SQL
@@
In the ToT baseline, the agent receives the entire database schema at once and is asked to generate the complete SQL query that satisfies the NLQ. This configuration represents global reasoning over the full relational structure, without any decomposition or locality constraints.
@@
B.5 Implementation
We implement Spider experiment as a Terminal-Bench Team (2025) task to ensure reproducibility and comparability between agents.
Appendix C Additional Analysis: Cost Evaluation
To assess more different methods and the token costs, we randomly selected tasks from SAT-Bench covering the joint distribution of treewidth and bag count induced by tree decomposition and tested with Chain of Thought (CoT), Divide and Conquer (DaC), Tree of Thoughts (ToT), and ACONIC(Ours). Specifically, we uniformly sample problem instances across discrete bins of $(\mathrm{tw}(G),|T|)$, where $\mathrm{tw}(G)$ denotes the treewidth of the constraint graph and $|T|$ the number of bags in the corresponding decomposition. This ensures that the evaluation set is not biased toward either very simple or very complex structural regimes.
Table 1 reports the aggregate accuracy and average token usage of each method under this sampling scheme.
Method Accuracy Avg. Tokens
CoT 21.95% 24,179
DaC 21.95% 129,201
ToT 26.82% 209,976
ACONIC 29.27% 19,247
Table 1: Performance under random structural coverage across all treewidth and bag-count distributions.
ACONIC achieves the highest accuracy while using fewer tokens than all baselines. This suggests that explicitly exploiting structural decompositions yields both effectiveness and efficiency benefits when reasoning over constraint-based tasks.
We also observe that Divide-and-Conquer (DaC) performs comparably to chain-of-thought in accuracy but incurs substantially higher token cost. This behavior is consistent with the nature of SAT and Spider tasks, which require maintaining and propagating multiple dependent variable assignments across steps. In such settings, naively decomposing the task into independent subquestions can lead to excessive context expansion without a corresponding gain in correctness.
Overall, these results support the view that structure-aware decomposition guided by the underlying constraint graph is better aligned with tasks characterized by strong variable dependencies, whereas generic divide-and-conquer strategies are less effective when subproblems are not weakly coupled.
Appendix D Additional Analysis: Tree of Thoughts
We include results for Tree of Thoughts (ToT) on SAT-Bench (See figure 6. Under our ToT configuration, ToT solves 34.8% of tasks, compared to CoT’s 21.5% and ACONIC’s 36.5%.
Method CoT ToT ACONIC
Success Rate 21.5% 34.8% 36.5%
Table 2: SAT-Bench success rate comparison across reasoning methods.
(a) Llama Chain of Thoughts
(b) Llama Aconic ACONIC
(c) Llama Tree of Thoughts
Figure 6: Task difficulty frontiers for LLaMA under CoT, ACONIC, and ToT on SAT-Bench. Each subplot shows success rate distribution based on problem complexity.
We follow the implementation in the original paper Yao et al. (2023). The search procedure is parameterized by: (i) maximum depth T=5T{=}5, i.e., the search tree expands for at most 5 levels; (ii) branching factor b=3b{=}3, i.e., each node proposes 3 candidate thoughts; and (iii) frontier size k=5k{=}5, i.e., we keep the top 5 candidates at each depth as the frontier for further expansion.
Appendix E Boolean Encoding and SAT-to-CSP Reduction (Meeting Scheduling Example)
We provide the mechanical steps that instantiate the meeting scheduling example as a Boolean formula and then as a CSP.
Variables and domains.
Let $\mathcal{P}$ denote the set of participants, $\mathcal{T}$ the set of time slots, and $\mathcal{L}$ the set of locations. For each person $p\in\mathcal{P}$ and slot $(t,\ell)\in\mathcal{T}\times\mathcal{L}$, introduce a Boolean variable
$x_{p}^{(t,\ell)}\in\{0,1\},$
where $x_{p}^{(t,\ell)}=1$ indicates that $p$ attends slot $(t,\ell)$. For each person $p$, let $\mathcal{A}_{p}\subseteq\mathcal{T}\times\mathcal{L}$ be the feasible availability pairs extracted from the system record. For each desired meeting pair $(p,p^{\prime})$, define the pairwise feasible set
$\mathcal{F}_{p,p^{\prime}}:=\mathcal{A}_{p}\cap\mathcal{A}_{p^{\prime}}.$
Boolean formulation of constraints.
Define meeting indicators for a pair $(p,p^{\prime})$:
$M_{p,p^{\prime}}^{(t,\ell)}:=x_{p}^{(t,\ell)}\wedge x_{p^{\prime}}^{(t,\ell)}.$
The task constraints consist of: (i) Exactly-one feasible meeting slot for each desired pair, and (ii) No simultaneous meetings for a person. Concretely,
$\bigwedge_{(p,p^{\prime})\in\{(A,B),(A,C)\}}\mathrm{EO}\Big(\{M_{p,p^{\prime}}^{(t,\ell)}\mid(t,\ell)\in\mathcal{F}_{p,p^{\prime}}\}\Big)$ (1)
$\wedge\ \bigwedge_{(t,\ell)}\neg\Big(M_{A,B}^{(t,\ell)}\wedge M_{A,C}^{(t,\ell)}\Big),$ (2)
where $\mathrm{EO}$ denotes Exactly One.
SAT (CNF) representation.
Let $\Phi=\{C_{1},\dots,C_{m}\}$ denote the resulting CNF clauses, where each clause is a disjunction of literals over Boolean atoms (e.g., $x_{p}^{(t,\ell)}$ or $\neg x_{p}^{(t,\ell)}$). The meeting scheduling instance is satisfiable iff there exists an assignment to all involved Boolean atoms that satisfies every clause in $\Phi$.
Reduction to CSP.
Given $\Phi$, we construct a CSP $\mathcal{R}=\langle X,D,\mathcal{C}\rangle$ as follows.
Variables. $X$ contains one CSP variable for each Boolean atom appearing in $\Phi$. For notational simplicity, we reuse the same symbol for the CSP variable (e.g., $x_{p}^{(t,\ell)}$).
Domains. Each variable is Boolean with domain
$D_{x}=\{0,1\}.$
Constraints. For each CNF clause $C_{j}$ with variable set $\mathrm{vars}(C_{j})\subseteq X$, we introduce $\mathbf{T}_{j}$ where $\mathbf{T}_{j}$ contains exactly those local assignments to $\mathrm{vars}(C_{j})$ that satisfy clause $C_{j}$. Letting $\mathcal{C}=\{\mathbf{T}_{1},\dots,\mathbf{T}_{m}\}$, an assignment satisfies the CSP iff it satisfies all clause relations.
Equivalence.
By construction, any satisfying assignment of the CNF formula $\Phi$ induces a satisfying assignment of $\mathcal{C}$, and any satisfying assignment of $\mathcal{C}$ satisfies every clause in $\Phi$. Therefore, solving the resulting CSP is equivalent to solving the original meeting scheduling instance.